An insurance sales representative sells policies to 5 men, all identical age and in good health.
According to actual tables, the probability that a man of this particular age will be alive 30
years hence is 2/3. Find the probability that in 30 years
(a) all 5 men; (b) at least 3 men; (c) none will be alive
1
Expert's answer
2019-12-18T11:19:54-0500

Let "X=" the number of men, who will be alive in 30 years: "X\sim B(n, p)"


"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}"

Given that "n=5, p={2 \over 3}"

(a)


"P(X=5)=\binom{5}{5}({2 \over 3})^5(1-{2 \over 3})^{5-5}={32 \over 243}\approx0.1317"

(b)


"P(X\geq3)=P(X=3)+P(X=4)+P(X=5)="

"=\binom{5}{3}({2 \over 3})^3(1-{2 \over 3})^{5-3}+\binom{5}{4}({2 \over 3})^4(1-{2 \over 3})^{5-4}+"

"+\binom{5}{5}({2 \over 3})^5(1-{2 \over 3})^{5-5}={80 \over 243}+{80 \over 243}+{32 \over 243}="

"={64 \over 81}\approx0.7901"

(c)


"P(X=0)=\binom{5}{0}({2 \over 3})^0(1-{2 \over 3})^{5-0}={1 \over 243}\approx0.0041"